If X and Y are independent random variables and if x has variable 5 and y has 2 then the variance of 3x+2y is ?
Answer by stanbon(75887)

If X and Y are independent random variables and if x has variable 5 and y has 2 then the variance of 3x+2y is ?
------------
That is a bit vague.
3x+2y = 3*5+2*2 = 15+4 = 19 ?
=================================
Cheers,
Stan H